What is the volume of a right pyramid whose base is a square with a side 6 m long and whose altitude is equal to the base side?
A. 36 m3
B. 72 m3
C. 108 m3
D. 216 m3

Answer by stanbon(75887)

What is the volume of a right pyramid whose base is a square with a side 6 m long and whose altitude is equal to the base side?
------
Volume = (1/3)(base area)(height)
----
V = (1/3)*6^2*6 = 72 m^3
======================
Cheers,
Stan
-------------------
A. 36 m3
B. 72 m3
C. 108 m3
D. 216 m3